# Demystifying CLIP Data

### v5

@@ 1 Introduction
Figure 1: ViT-B/32 on ImageNet zero-shot classification with fixed training steps (12.8B seen pairs and training/validation data has been de-duplicated). Raw: raw CommonCrawl (CC) distribution; Raw English: English only CC; MetaCLIP w/o bal.: curated (sub-string matched) data pool from CC; MetaCLIP: curated and balanced metadata distribution. Metadata curation boosts performance significantly and balancing is equally important. Our MetaCLIP data significantly outperforms CLIP’s WIT400M and LAION data(Schuhmann et al., 2021).
@@ 3 MetaCLIP
“To address this, we constructed a new dataset of 400 million (image, text) pairs collected from a variety of publicly available sources on the Internet. To attempt to cover as broad a set of visual concepts as possible, we search for (image, text) pairs as part of the construction process whose text includes one of a set of 500,000 queries We approximately class balance the results by including up to 20,000 (image, text) pairs per query.
@@
3.1 Metadata construction: =$\mathcal{M}=\{\textit{entry}\}$
@@
“The base query list is all words occurring at least 100 times in the English version of Wikipedia. This is augmented with bi-grams with high pointwise mutual information as well as the names of all Wikipedia articles above a certain search volume. Finally all WordNet synsets not already in the query list are added.
@@
[…]
@@
“We also restrict this step in CLIP to text-only querying for sub-string matches while most webly supervised work uses standard image search engines …
@@
[…]
@@ 3.4 Query and Balancing with $t\leq$20k
The magic number $t=20$k is a threshold used to limit the number of texts/pairs for each entry. Entries with fewer than $t$ pairs (tail entries) retain all associated pairs, while entries with more than $t$ pairs (head entries) are sub-sampled to $t$ pairs. The selection is based on the density of information in texts; texts with more matched entries have a higher chance of being curated (recall that the average is 3.5 matches per text).
To study the effect of the magic number $t=20$k, we plot the cumulative sum of counts for entries sorted by counts from tail to head in Fig. 2. Interestingly, the value of $t=20$k seemingly represents the transition from tail to head entries, when the head entries start exhibiting an exponential growth rate. By applying a max count of $t$, the growth rate of total counts (i.e., the scale of resulting data points) is reduced to linear. This significantly flattens (and balances) the training data distribution. We further study the optimality of $t=20$k for the 400M data scale in our experiments.
Figure 2: Cumulative sum of counts on entries from tail to head on a data pool with 1.6B image-text pairs (5.6B match counts). (1) raw/unbalanced cumulative counts, $t=\infty$; (2) balanced cumulative counts after applying $t=20$k. The limit $t$ defines the transition of tail/head entries.
@@
(i) It reduces dominance and noise from head entries, like common web terms. E.g., out of 400M pairs, only $20$k texts containing “photo” are kept (while there are 54M “photo” instances in the pool).
@@ 3.5 A simple Algorithm for Curation
[…]
@@ 4.1 Results
Table 5: Scaling MetaCLIP from 400M ($t$=20k) to 1B ($t$=20k) and 2.5B ($t$=170k) training data.
@@
Figure 3: Cumulative sum of counts on entries from tail to head on a data Pool 2. We again show (1) raw/unbalanced cumulative counts), $t=\infty$; (2) balanced cumulative counts after applying $t=20$k and $t=170$k. $t$ defines maximum number of counts per entry and the transition of tail/head entries. We show the Pool 1 configuration from Fig. 2 as dashed lines for reference.
@@ 4.2 Ablation Study
To understand the key effect of balancing, we use the whole matched pool (1.6B image-text pairs) to train CLIP. Surprisingly, training on 4$\times$ more data (on head entries) significantly hurts the accuracy on ImageNet (61.9 vs 65.5) and averaged accuracy across 26 tasks (56.6 vs 58.2).
@@
Table 6: Ablation studies on balancing in MetaCLIP. Default: $t$=20k, 400M. Model: ViT-B/32.
@@ A.1 Additional Results
The concurrent work Gadre et al. (2023) released a collection of 12.8B image-text pairs from CommonCrawl from 2014-2022. We further investigate whether we can apply the algorithm on its 12.8 unfiltered pool. Although the unfiltered pool seemingly offers an opportunity to apply our algorithm on a publicly available source, our initial studies show that, implicit biases may still be present in this pool. For example, we notice that all image URLs are collected as a string starting with http. This excludes relative URLs that could be frequently used by quality websites (with potentially good image-text pairs). We curate from DataComp’s 12.8B unfiltered pool with $t$=60k, which has 6% of tail counts that is the same as $t$=20k for 400M from our 1.6B pool.
@@ A.2 Details on Efficient Curation
Specifically, sub-string matching can be placed immediately after HTML parsing to reduce data points for English-only pairs (e.g. by $\sim$50%). Balancing can be applied earlier, before image downloading, to further reduce data points by $\sim$77%. This approach led to a total reduction of $\sim$90%, allowing for curation of the entire CommonCrawl data without the need to store and transfer all data points, which allows us to curate the whole CommonCrawl since 2013 with 300B+ scale URL-text pairs without spending the storage/transfer on all $\sim$10$\times$ data points, where the rate of keeping a data point with MetaCLIP curation is $\sim$0.1 (0.5 $\times$ 0.23).
@@ A.3 Human Study on the Effects of Curation
Image $4.60$, $[4.50,4.70]$ $4.36$, $[4.23,4.48]$ $<0.001$
Text $4.67$, $[4.56,4.78]$ $4.06$, $[3.82,4.30]$ $<0.001$
Alignment $4.41$, $[4.23,4.59]$ $3.72$, $[3.46,3.99]$ $<0.001$
@@ A.5 Training Setup of OpenAI CLIP vs OpenCLIP
Our work strictly follows CLIP’s setup for a controlled comparison focusing on data curation and quality. We notice differences in the training setup of OpenCLIP [^5] https://github.com/mlfoundations/open_clip and list the difference (known to us). OpenCLIP varies the setup from CLIP (e.g., global batch size, learning schedule, etc.). Here we only list the difference for LAION-400M(Schuhmann et al., 2021), which is closer to the CLIP setup. We note that DataComp differs even more, e.g., by curating images close to ImageNet training data, a large batch size of 90k that is almost 3$\times$ larger than CLIP, and using the CLIP model to filter data.
@@
Seen Pairs 12.8B(400M×$\times$32 epochs) 13B (407M×$\times$32 epochs) 12.8B
@@ A.10 Qualitative Data Examples
Contributions
Hu Xu: project lead, overall direction, data pipeline (parsing, dedup, download and mitigation), training, efficiency optimization, evaluation and writing.
Saining Xie: advising on directions, impact, writing and define key terms.
Xiaoqing Ellen Tan: data ablation, tuning mitigation and performance tradeoff.
Po-Yao Huang: DataComp benchmark evaluation, porting code to new cluster.
Russell Howes: advising on adjusting data pipeline, suggestions on scaling of URLs deduplication.
Vasu Sharma: running face blurring on 2.5B collection.
Shang-Wen Li: team management, overall suggestion on direction, data collection for 2 rounds, data mitigation, legal communication, data transfer, GPU allocation, coordination and public communication.
Gargi Ghosh: team management, GPU allocation.
Luke Zettlemoyer: general advising on direction and writing.
Christoph Feichtenhofer: advising on experimental setups, multi-round writing and editing.

### v6

@@ 1 Introduction
Figure 1: ViT-B/32 on ImageNet zero-shot classification with fixed training steps (12.8B seen pairs and training/validation data has been de-duplicated). Raw: raw CommonCrawl (CC) distribution; Raw English: English only $C$; MetaCLIP w/o bal.: curated (sub-string matched) data pool from $C$; MetaCLIP: curated and balanced metadata distribution. Metadata curation boosts performance significantly and balancing is equally important. Our MetaCLIP data significantly outperforms CLIP’s WIT400M and LAION data(Schuhmann et al., 2021).
@@ 3 MetaCLIP
“To address this, we constructed a new dataset of 400 million (image, text) pairs collected from a variety of publicly available sources on the Internet. To attempt to cover as broad a set of visual concepts as possible, we search for (image, text) pairs as part of the construction process whose text includes one of a set of 500,000 queries We approximately class balance the results by including up to 20,000 (image, text) pairs per query.   ”
@@
3.1 Metadata construction: $\mathcal{M}=\{\textit{entry}\}$
@@
“The base query list is all words occurring at least 100 times in the English version of Wikipedia. This is augmented with bi-grams with high pointwise mutual information as well as the names of all Wikipedia articles above a certain search volume. Finally all WordNet synsets not already in the query list are added.   ”
@@
3.2 Sub-string Matching: $\textit{text}\rightarrow\textit{entry}$
@@
“We also restrict this step in CLIP to text-only querying for sub-string matches while most webly supervised work uses standard image search engines …   ”
@@
As a result, sub-string matching builds the mapping $\textit{txt}\rightarrow\textit{entry}$. This step has two outcomes: (1) low-quality text is dropped; (2) unstructured text now has a structured association with metadata. For all English text, $\sim$50% image-text pairs are kept in this stage. Similar to CiT (Xu et al., 2023), this approach looks for quality matches and automatically gets rid of some type of noise (such as date strings) that a typical filter system would remove consider case-by-case (e.g., regular expression on dates, ids etc.).
3.3 Inverted Indexing: $\textit{entry}\rightarrow\textit{text}$
Following sub-string matching, CLIP builds an inverted index of the data pool. All texts associated with each metadata entry are aggregated into lists, creating a mapping from each entry to the corresponding texts, $\textit{entry}\rightarrow\textit{text}$.
@@ 3.4 Query and Balancing with $t\leq$20k
The magic number t=20t=20k is a threshold used to limit the number of texts/pairs for each entry. Entries with fewer than $t$ pairs (tail entries) retain all associated pairs, while entries with more than $t$ pairs (head entries) are sub-sampled to $t$ pairs. The selection is based on the density of information in texts; texts with more matched entries have a higher chance of being curated (recall that the average is 3.5 matches per text).
To study the effect of the magic number t=20t=20k, we plot the cumulative sum of counts for entries sorted by counts from tail to head in Fig. 2. Interestingly, the value of t=20t=20k seemingly represents the transition from tail to head entries, when the head entries start exhibiting an exponential growth rate. By applying a max count of $t$, the growth rate of total counts (i.e., the scale of resulting data points) is reduced to linear. This significantly flattens (and balances) the training data distribution. We further study the optimality of t=20t=20k for the 400M data scale in our experiments.
Figure 2: Cumulative sum of counts on entries from tail to head on a data pool with 1.6B image-text pairs (5.6B match counts). (1) raw/unbalanced cumulative counts, $t=\infty$; (2) balanced cumulative counts after applying t=20t=20k. The limit $t$ defines the transition of tail/head entries.
@@
(i) It reduces dominance and noise from head entries, like common web terms. E.g., out of 400M pairs, only 2020k texts containing “photo” are kept (while there are 54M “photo” instances in the pool).
@@ 3.5 A simple Algorithm for Curation
This procedure is equivalent to CLIP’s curation, because if one image-text pair has one or more matched entries, the chance of that pair being selected is determined by the probability of sampling for each individual entry: $t/\texttt{entry\_count[entry\_id]}$. As long as one entry selects that pair, it will be kept in $\mathcal{D}^{*}$. Our independent sampling approach allows us to scale balancing for each data point independently and reduces the global operation to counting the total matches for each entry. We demonstrate case studies in experiments on (1) scaling curation in a data pipeline and (2) online balancing in data loader.
@@ 4.1 Results
Table 5: Scaling MetaCLIP from 400M (tt=20k) to 1B (tt=20k) and 2.5B (tt=170k) training data.
@@
Figure 3: Cumulative sum of counts on entries from tail to head on a data Pool 2. We again show (1) raw/unbalanced cumulative counts), $t=\infty$; (2) balanced cumulative counts after applying t=20t=20k and t=170t=170k. $t$ defines maximum number of counts per entry and the transition of tail/head entries. We show the Pool 1 configuration from Fig. 2 as dashed lines for reference.
@@ 4.2 Ablation Study
[…]
@@
Table 6: Ablation studies on balancing in MetaCLIP. Default: tt=20k, 400M. Model: ViT-B/32.
@@ A.1 Additional Results
The concurrent work Gadre et al. (2023) released a collection of 12.8B image-text pairs from CommonCrawl from 2014-2022. We further investigate whether we can apply the algorithm on its 12.8 unfiltered pool. Although the unfiltered pool seemingly offers an opportunity to apply our algorithm on a publicly available source, our initial studies show that, implicit biases may still be present in this pool. For example, we notice that all image URLs are collected as a string starting with http. This excludes relative URLs that could be frequently used by quality websites (with potentially good image-text pairs). We curate from DataComp’s 12.8B unfiltered pool with tt=60k, which has 6% of tail counts that is the same as tt=20k for 400M from our 1.6B pool.
@@ A.2 Details on Efficient Curation
[…]
@@ A.3 Human Study on the Effects of Curation
Image $4.60$, $[4.50,4.70]$ $4.36$, $[4.23,4.48]$ <0.001<0.001
Text $4.67$, $[4.56,4.78]$ $4.06$, $[3.82,4.30]$ <0.001<0.001
Alignment $4.41$, $[4.23,4.59]$ $3.72$, $[3.46,3.99]$ <0.001<0.001
@@ A.5 Training Setup of OpenAI CLIP vs OpenCLIP
[…]
@@
[…]